OpenStudy (anonymous):

which value for x makes th esentence truw? 6x=5x+5

OpenStudy (anonymous):

do you know how to solve for x

OpenStudy (anonymous):

Well, for this one you could apply mental math. For instance, see if you can use the number 5 . . . if 5 were the value of x... 6(5) = 5(5) +5 Which would be . . . 30=25+5 30=30. The value of x being 5 makes the equation true and because you worked it out, like I did both mentally and by plugging it into the original problem.

OpenStudy (anonymous):

then the answer would be five

OpenStudy (anonymous):

Precisely.

OpenStudy (anonymous):

you didnt need help with this you just answerd your own question lol